A company expected its annual overhead costs to be $20,000 and direct labor hours to be 10,000. Actual overhead was $25,000, and actual labor hours totaled 12,000. Calculate predetermined overhead rate, amount of applied overhead and state whether it results in under applied or over applied.
Answer #1

Predetermined Overhead rate

= Estimated overhead/Estimated direct labor hours

= 20,000/10,000

= $2 per direct labor hour

.

Applied overhead rate

= Predetermined overhead rate * Actual direct labor hours

= 2 * 12,000

= $24,000

.

Actual overhead = $25,000

Applied overhead = $24,000

Overhead underapplied = 25,000 - 24,000

= $1,000 underapplied